Fast Turnstiles Co. is evaluating the extension of credit to a new group of customers. Although these customers will provide $108,000 in additional credit sales, 8 percent are likely to be uncollectible. The company will also incur $15,800 in additional collection expense. Production and marketing costs represent 71 percent of sales. The firm is in a 30 percent tax bracket and has a receivables turnover of four times. No other asset buildup will be required to service the new customers. The firm has a 8 percent desired return.  

a-1. Calculate the incremental income after taxes.  
  

    

a-2. Calculate the return on incremental investment. (Input your answer as a percent rounded to 2 decimal places.)
  

   

a-3. Should Fast Turnstiles Co. extend credit to these customers?
  

Yes
No

  

b-1. Calculate the incremental income after taxes if 11 percent of the new sales prove to be uncollectible.
  

   

b-2. Calculate the return on incremental investment if 11 percent of the new sales prove to be uncollectible. (Input your answer as a percent rounded to 2 decimal places.)
  

     

b-3. Should credit be extended if 11 percent of the new sales prove uncollectible?
  

Yes
No

  

c-1. Calculate the return on incremental investment if the receivables turnover drops to 1.6, and 8 percent of the accounts are uncollectible. (Input your answer as a percent rounded to 2 decimal places.)
  

   

c-2. Should credit be extended if the receivables turnover drops to 1.6, and 8 percent of the accounts are uncollectible?
  

Yes
No
Answer #1

a-1. Incremental income after tax will be as followed :-

Sales 108000
(-) production and marketing cost (71% of 108000) (76680)
(-) additional collection exp (15800)
(-) loss due to non collectability (8% of 108000) (8640)
Income before tax $6880
(-) tax expense (30% of 6880) (2064)
Income after tax $4816

a-2. Incremental investment will be investment in working capital which will be accounts receivable .

We have receivable turnover = 4

That is to say, sales / accounts receivable = 4

Thus, accounts receivable = 108000/4 = $27000 in new receivables

Return on incremental investment = income after tax / incremental investment x100

Therefore return on incremental investment = 4816/27000x100 = 17.84%.

a-3 . Desired rate of return of Fast turnstiles co. is 8%. Whereas it is able to earn 17.84% on it's incremental investment . Since the company is earning more than what is required, therefore it should extend the credit to the customer.

Therefore correct answer is Yes.

b-1 . If 11% of the sales is uncollectible then income after tax will be as follows :-

Sales 108000
(-) production and marketing exp (71% of 108000) (76680)
(-) collection exp (15800)
(-) loss due to non collection (11% of 108000) (11880)
Income before tax $3640
(-) tax (30% of 3640) (1092)
Income after tax $2548

b-2 . Return on incremental investment if 11 % of sales is uncollectible = 2548/27000x100 = 9.44%.

b-3 . Desired return of the firm is 8% whereas return from extending credit to new customer is coming out to be 9.44%. since firm is getting more return than desired if it extend credit. Therefore , firm should extend credit.

Correct answer is Yes.

c-1. If assets turnover drops to 1.6, then new additional investment will be

108000/1.6 = $67500.

Income after tax if uncollectible rate is 8% = $4816.

Therefore return on incremental investment = 4816/67500x100 = 7.13%.

c-2 . Desired return is 8% whereas we are getting Return on incremental investment as 7.13%. since it is less than desired return therefore we firm should avoid extending credit.

Correct answer is No.